OpenStudy (anonymous):

Graph the function f(x)=-abs(2x-3). Be sure to find any x and y intercepts, places where the graph has a relative maximum or minimum.

OpenStudy (anonymous):

2x -3 = 0 2x = 3 x=1.5 now if x>1.5 the expression is positive and f(x) = -(2x-3) = 3 -2x if x<1.5 the expression is negative and f(x) = 2x - 3 if x = 1.5 f(x) = 0